OpenStudy (anonymous):

Evaluate each expression for the given value(s) of the variable(s) (m-4)-5 (<---Thats to the negative fifth power, that -5) for m=6

OpenStudy (anonymous):

\[(m-4)^{-5}=\frac{1}{(m-4)^5}\]

OpenStudy (anonymous):

Thank you! :D wow that was way more simple than I thought..

OpenStudy (anonymous):

if \[m=6\] you get \[\frac{1}{(6-4)^5}=\frac{1}{2^5}=\frac{1}{32}\]

OpenStudy (anonymous):

yw